OpenStudy (anonymous):

What percent of 20 is 18. Plz show ur work.

OpenStudy (mathstudent55):

Use a proportion. 20 is 100%. 18 is what percent? 20/100 = 18/x Cross multiply: 20x = 100(18) 20x = 1800 x = 90

OpenStudy (anonymous):

thanx alot

OpenStudy (mathstudent55):

wlcm